consider a cylinder of a square section 10cm side which strouhal number is about 0.13 it has mass of 1 kgand mounted on a spring of stiffness 1 N/m give value of flow velocity at which viv occur in m/s
1
Expert's answer
2020-07-21T12:38:06-0400

"Sr=\frac{f\cdot l}{v}"


"f=\frac{1}{2\pi\sqrt{m/k}}=\frac{1}{2\pi\sqrt{1/1}}=0.16Hz"


"v=\frac{f\cdot l}{Sr}=\frac{0.16\cdot 0.1}{0.13}=0.12m/s"